The cost of winning a seat in the House in recent years has been approximately 805e^0.0625x thousand dollars, where x is the number of years since 2000. Estimate the cost of winning a House seat in the year 2010.: The cost of winning a seat in the House in recent years has been approximately 805e^0.0625x thousand dollars, where x is the number of years since 2000. Estimate the cost of winning a House seat in the year 2010.
Answer by ankor@dixie-net.com(4489)

C(x) = 805e^0.0625x thousand dollars,
where x is the number of years since 2000.
:
Estimate the cost of winning a House seat in the year 2010.
x = 10
:
C(x) = 805(e^(.0625x10))
:
C(x) = 805(e^.625)
:
Find e^.625 on a good calc
c(x) = 805 * 1.868246
:
C(x) = 1504 (thousand$)